The daily revenue of Modern Style Hair Salon is approximately normally distributed with a mean of R7200 and a standard deviation of R1500. 75% of days, the salon makes more than (Rands only no cents)


1
Expert's answer
2021-11-15T20:03:09-0500

"P(Z>z)=0.75\to P(Z<z)=0.25)\to z=-0.67."

"\frac{x-\mu}{\sigma}=-0.67\to x=\mu-0.67\sigma=7200-0.67*1500=6195."